# Semiparametric Skew-Elliptical Distributions For High-Dimensional Graphical Models

### v1

@@ 3 The elliptical skew-(S)KEPTIC estimator
where $D=\sqrt{\frac{2}{\pi}}D_{\alpha}(I_{p}+D^{2}_{\alpha})^{-\frac{1}{2}}$.
Lemma 3.4.
@@
$\boldsymbol{X}\sim\operatorname{SU}\mathcal{E}_{p,p}(\xi,\Sigma,\Lambda,h^{(2p)},0,\Gamma)$ reduces to a $\operatorname{SUT}$ distribution: $\boldsymbol{X}\sim\operatorname{SUT}_{p,p}(\xi,\Sigma,\Lambda,\nu,0,\Gamma)$.
Theorem 3.5 (elliptical skew-KEPTIC estimator).
Assuming $\boldsymbol{X}\sim MS\mathcal{E}_{p,p}(\Sigma^{0},0,\Lambda;f)$, such that $\boldsymbol{Z}=f(\boldsymbol{X})$ denotes a generic $\operatorname{SU\mathcal{E}}$ distribution: $\boldsymbol{Z}\sim\operatorname{SU\mathcal{E}}_{p,p}(0,\Sigma^{0},\Lambda,h^{(2p)},0,\Gamma)$. The elliptical skew-KEPTIC estimator is:
As discussed in Hult and Lindskog, (2002), Spearman’s rho is not invariant in the class of elliptical distributions, and its relation with linear correlation coefficient in Lemma 2.1 holds if and only if the distribution is nonparanonrmal.
Proof.
We consider $\boldsymbol{X}\sim MS\mathcal{E}_{p,p}(\Sigma^{0},0,\Lambda;f)$, where $f(\boldsymbol{X})\sim V_{0}^{-1/2}Z_{0}$, and $Z_{0}\sim\operatorname{CSN}_{p,p}(0,\Omega^{-1},D_{\alpha}L,0,D_{\kappa})$ where $D_{\kappa}=I_{p}$, and $V_{0}=1$. We compute $\mathbb{C}orr(X)$ using Kendall’s tau statistic, $\Sigma^{0}_{\tau}=\left[\tau_{ij}\right]$, and we decompose $\Sigma^{0}_{\tau}=L^{\top}D_{\kappa}L$, aligning it $\Sigma^{0}_{\tau}$ to the assumption of $\boldsymbol{X}\sim MS\mathcal{E}_{p,p}(\Sigma^{0},0,\Lambda;f)$ exploiting (9).
[…]
$=(L^{\top}ML)^{-1}\quad\text{where:}\quad M=I-D^{2},$
$=\frac{Adj(L^{\top}ML)}{det(L^{\top}ML)},$
$=\frac{Adj(L^{\top}ML)}{det(L)^{2}det(M)}.$
Where $Adj$ is the adjoint matrix, and $det$ is the determinant of the matrix. The matrices $I$ and $D^{2}$ are both diagonal matrices, and the matrix $M$ is. Thus, we have a new decomposition of the correlation matrix in which the skewness parameter $\alpha$, belonging to the matrix $D_{\alpha}$, contributes to the determination of the linear relations. The matrix $(L^{\top}ML)$ is:
where the generic $m_{k,k}$ term is:
$m_{kk}=I_{kk}-D^{2}_{kk}=1-\frac{2}{\pi}\frac{\alpha_{kk}^{2}}{1+\alpha_{kk}^{2}}.$
Let $A_{{\tau}}$ denote the final correlation matrix of $X$ with Kendall’s tau statistic. We recall that the stochastic representation (8) of the $\operatorname{CSN}$ is a sum of two comonotic copula functions that have the invariance property, so the SKEPTIC estimator is the same for both the distributions, thus the elliptical skew-KEPTIC estimator is the sum of two SKEPTIC estimators and an estimated matrix, $A_{\tau}$, from $f(\boldsymbol{X})\sim\operatorname{CSN}_{p,p}(0,\Omega^{-1},D_{\alpha}L,0,D_{\kappa})$.
When $V_{0}\sim Gamma(\nu/2,\nu/2)$, we have that $f(\boldsymbol{X})\sim\operatorname{SUT}_{p,p}(\xi,\Sigma,\Lambda,\nu,0,\Gamma)$ or equivalently: $f(\boldsymbol{X})\sim\frac{Z_{0}}{\sqrt{V_{0}}}$. Since the positive support of Gamma random variable, and the fact that Kendall’s tau is invariant under monotone transformations, we state that the skew-KEPTIC estimator maintains the same ordinal measures of dependencies of CSN distribution even when $f(\boldsymbol{X})\sim\operatorname{SUT}_{p,p}(0,\Sigma^{0},\Lambda,\nu,0,\Gamma)$, where $\Gamma=I_{p}$. ∎
Lemma 3.6.
(skew-Normal SKEPTIC estimator) Assuming $\boldsymbol{X}\sim MS\mathcal{E}_{p,p}(\Sigma^{0},0,\Lambda;f)$, such that $\boldsymbol{Z}=f(\boldsymbol{X})$ denotes a generic $\operatorname{SU\mathcal{E}}$ distribution, when $\boldsymbol{Z}\sim\operatorname{SU}\mathcal{E}_{p,p}(0,\Sigma^{0},\Lambda,h^{(2p)},0,\Gamma)$ reduces to a $\operatorname{CSN}$ distribution, the skew-Normal SKEPTIC estimator is:
Proof.
We recall Lemma 2.1, where for the Normal distribution, and for $i\neq j$, we have: ${S}_{ij}=2\sin\left(\frac{\pi}{6}{\rho}_{ij}\right)=\sin\left(\frac{\pi}{2}{\tau}_{ij}\right)$. Let $A_{\rho_{ij}}$ denote the final correlation matrix of $\boldsymbol{X}$ with Spearman’s rho statistic, as in the Normal distribution, we have: ${S}^{\rho,\alpha}_{ij}\equiv 4\sin\left(\frac{\pi}{6}A_{{\rho}_{ij}}\right)=2\sin\left(\frac{\pi}{2}A_{{\tau}_{ij}}\right)$. ∎
We generally refer to the estimators in Theorem 3.5 and Lemma 3.6 as elliptical skew-(S)KEPTIC estimators. The specification of the inclusion of the Spearman’s rho will be clear in the application. To obtain the correlation matrix using the elliptical skew-(S)KEPTIC estimator, we need to estimate the parameters of the $\operatorname{CSN}$ distribution. The estimation procedure is based on the EM algorithm of Abe et al., (2021) [^1] The estimated parameters come from the skew-Normal distribution of Azzalini and Capitanio, (1999), with an adequate parametrization explained in Arellano-Valle and Azzalini, (2006) we conform the parameters to a CSN distribution..
3.1 Estimating Marginal Distributions
We find estimators $\{f_{j}\}_{j=1}^{p}$ such that the transformed data are distributed according to a $\operatorname{CSN}$. Let $\Psi$ denote the cumulative distribution function of the $\operatorname{CSN}$ distribution, then the estimators are:
$\hat{f}_{csn,j}(x_{ij})=\Psi^{-1}(\hat{r}_{ij}),\quad\text{where: }\hat{r}_{ij}=\frac{r_{ij}}{n+1}.$ (11)
The functions $\{f_{csn,j}(x_{ij})\}_{j=1}^{p}$ are monotone and differentiable, and Gupta et al., (2004) show that $\operatorname{CSN}$ distribution is closed under conditioning therefore, according to Sklar’s theorem, (6) is a $\operatorname{CSN}$ copula.
3.2 Methods for the Estimation of the Precision Matrix and the Undirected Graph
Upon completing the estimation of the correlation matrix, we follow Liu et al., 2012a and estimate the precision matrix capturing the interrelationships among variables within an undirected graph.
3.2.1 The Dantzig selector
The Dantzig selector estimator (Yuan, , 2010) exploits the relation between the coefficients $\theta$ stemming from a multivariate linear regression and the elements of the precision matrix. Sparse precision matrix implies sparse regression coefficients estimation that relies on an algorithm that depends on the tuning parameter $\delta$. Dantzig procedure involves two steps:
  • Estimation: for $j=1,\ldots,p$, calculate
    $\hat{\theta}^{j}$ $=\text{arg}\min_{\theta\in\mathbb{R}^{p-1}}\|\theta\|_{1}\quad\text{subject to }\quad\|\hat{S}_{\backslash j,j}-\hat{S}_{\backslash j,\backslash j}\theta\|_{\infty}\leq\delta,$ (12)
    $\hat{\Omega}_{jj}$ $=\left[1-2(\hat{\theta}^{j})\hat{S}_{\backslash j,j}+(\hat{\theta}^{j})^{\top}\hat{S}_{\backslash j,\backslash j}\hat{\theta}^{j}\right]\quad\text{and }\quad\hat{\Omega}_{\backslash j.j}=-\hat{\Omega}_{\backslash jj}\hat{\omega}^{j}.$
@@ 3.2.2 The Glasso Algorithm
The penalty term, $\lambda_{G}$, is the parameter that controls the sparsity of the estimated matrix, i.e., the number of zeros in the precision matrix. Banerjee et al., (2008) shows that the shrinking procedure based on the tuning parameter $\lambda_{G}$ is directly proportional to the input covariance matrix. This means that greater values of $\lambda_{G}$ imply a greater presence of zeros in the estimated $\Omega$, therefore a minor number of interconnections are unveiled.
3.2.3 The CLIME Algorithm
CLIME stands for constrained $\ell_{1}$-minimization for inverse matrix estimation. This algorithm, proposed by Cai et al., (2011), is very attractive for high-dimensional data since the estimator is obtained by combining vector solutions solving a linear program. Let $\Omega$ denote a $p\times p$ precision matrix and $\{\hat{\Omega}_{1}\}$ be the solution set of the optimization problem:
$\text{min}\|\Omega\|_{1}\quad\text{subject to:}$: (15)
$\|\Sigma_{n}\Omega-I\|_{\infty}\leq\lambda_{C},$
where $\lambda_{C}$ is the tuning parameter. The final CLIME estimator of $\Omega_{0}$ is defined by the symmetrization of $\hat{\Omega}_{1}$:
$\hat{\Omega}=(\hat{\omega}_{ij}),\quad\text{where:}$: (16)
@@ 4 Theoretical properties
We analyze the theoretical properties of the elliptical skew-(S)KEPTIC estimator. The results suggest that our estimator is reliable when dealing with highly skewed data. As Liu et al., 2012a , our results show an exponential concentration rate in $\|\cdot\|_{\max}$. Moreover, we prove optimal bounds for shrinkage procedures. We show the results for the Dantzig selector, however, they can be easily extended to the CLIME and the Glasso procedures.
@@
Let $x_{1},...,x_{n}$ be $n$ observations from $\boldsymbol{X}\sim\operatorname{MS\mathcal{E}}_{p}(\Sigma^{0},\xi,\Lambda;f)$, restricted to the case $h^{(p)}=\phi^{(p)}(u)$, and let $\hat{S}^{\rho,\alpha}_{ij}=4\sin(\frac{\pi}{6}A_{\hat{\rho,}_{ij}})\operatorname{I}(i\neq j)+\operatorname{I}(i=j)$ denote the elliptical skew-SKEPTIC, then with probability at least $1-1/p^{2}$ we have:
$\|\hat{S}^{\rho,\alpha}_{ij}-\Sigma^{0}_{ij}\|_{\max}\leq 16\pi\sqrt{\frac{logp}{n}}.$ (18)
@@
Let $x_{1},...,x_{n}$ be $n$ observations from $\boldsymbol{X}\sim\operatorname{MS\mathcal{E}}_{p}(\Sigma^{0},\xi,\Lambda;f)$ and let $\hat{S}^{\tau,\alpha}_{ij}=2\sin(\frac{\pi}{2}A_{\hat{\tau}_{ij}})\operatorname{I}(i\neq j)+\operatorname{I}(i=j)$ denote the elliptical skew-(S)KEPTIC estimator, then with probability at least $1-1/p^{2}$ we have:
$\|\hat{S}^{\tau,\alpha}_{ij}-\Sigma^{0}_{ij}\|_{\max}\leq 4\pi\sqrt{\frac{logp}{n}}.$ (19)
Proofs of Theorems 4.1 and 4.2 are in Appendix A. The elliptical skew-(S)KEPTIC estimator achieves convergence in high-dimension, thus it is possible to define the following result.
@@
When we substitute the estimated matrices $\hat{S}^{\rho,\alpha}$ and $\hat{S}^{\tau,\alpha}$ into the parametric graphical lasso (or the CLIME or graphical Dantzig selector), the estimators provide a slower convergence than the Gaussian case ensuring consistency and graph recovery as a Gaussian model.
Proof.
The proof is based on the observation that the sample correlation matrix $\hat{S}$ is a sufficient statistic for all three shrinkage procedures: the graphical lasso, graphical Dantzig selector, and CLIME. As Yuan, (2010) points out, a sufficient condition to ensure that an estimator is consistent is if there exists some constant $c$, such that $\mathbb{P}\left(\|\hat{S}-\Sigma^{0}\|_{\max}>c\sqrt{\frac{\log p}{n}}\right)\leq 1-\frac{1}{p}$, which can be replaced by (18) and (19) in Theorems 4.1 and 4.2. ∎
The high-dimensional convergence of the elliptical skew-(S)KEPTIC estimator also means that the precision matrix is bounded. Similar results are found by Cai et al., (2011) and Yuan, (2010). In particular, let $\hat{\Omega}^{mse-s}$ denote the inverse correlation matrix using the elliptical skew-(S)KEPTIC estimator and consider the Dantzig selector as the shrinkage procedure to find $\hat{\Omega}$. We introduce a class of inverse matrix defined as $\mathcal{M}_{1}(\kappa,\tau,M):=\{\Omega:\Omega\succ 0,\text{diag}(\Omega^{-1})=1,\|\Omega\|_{1}\leq\kappa,\frac{1}{\tau}\leq\lambda_{\min}(\Omega)\leq\lambda_{\max}(\Omega)\leq\tau,\deg(\Omega)\leq M\}$, where $\kappa,\tau>1$. Given Theorem 4.3, we obtain the following result:
@@
For $1\leq q<\infty$, there exists a constant $c$ that depends on $\kappa$, $\tau$, $\lambda_{\min}(\Omega)$ and $\lambda_{\max}(\Omega)$, such that
$\sup_{\Omega\in\mathcal{M}_{1}(\kappa,\tau,M)}\|\hat{\Omega}^{mse-s}-\Omega\|_{q}=O_{P}\left(M\sqrt{\frac{\log p}{n}}\right).$ (20)
provided that $\lim_{n\rightarrow\infty}\frac{n}{M^{2}\log d}=\infty$ and $\delta=c\sqrt{\frac{\log d}{n}}$, for sufficiently large $c$, where $\delta$ is the parameter in (12).
Proof.
This result follows Yuan, (2010), which proves that the difference between the estimated precision matrix and the theoretical precision matrix is bounded by $\|\hat{S}-\Sigma^{0}\|_{\max}$, where $\hat{S}$ is the estimated correlation matrix and $\Sigma^{0}$ is the theoretical correlation matrix. In our case, to prove the bounds for the estimated correlation matrix we just replace $\hat{S}$ with $\hat{S}^{\rho,\alpha}$ and $\hat{S}^{\tau,\alpha}$ where the bounds are proved in Theorem 4.1 and 4.2. ∎
@@ 5 Experiments
We present the empirical performance of the elliptical skew-(S)KEPTIC methods on both synthetic and real datasets, comparing the proposed method throughout this paper with the SKEPTIC estimator in Liu et al., 2012a . In particular, we consider the following methods for the estimation of graphical models. We generate data from the Normal distribution, the CSN distribution, and the SUT distribution. For each generation, we compute the SKEPTIC and the SKEW-(S)KEPTIC estimator and conduct numerical simulations on synthetic data to show the ability of graph recovery of the proposed estimators. In conclusion, we present an empirical analysis where the methods are applied to the log-returns of the S&P500 index.
@@ 5.1 Numerical Simulations
The numerical simulations follow these steps. Using the R package huge (Jiang et al., (2021)), we generate the precision matrix $\Omega=(\Sigma^{0})^{-1}$ using the ”random” structure, in which we randomly generate pairs of conditional dependencies in the off-diagonal elements. We sample $n$ observations from $p$-variate Normal, $\operatorname{CSN}$ and $\operatorname{SUT}$ distribution. We set the location parameter equal to zero for the last two, and $\Sigma^{0}=\Omega^{-1}$. We consider the following cases: $p\in\{100,100,200\}$, and $n\in\{200,100,100\}$. The cases $n=100,p=100$, and $n=100,p=200$ consider the high-dimensional scenarios. Regarding the $\operatorname{CSN}$ parameters, we set $\Delta=1$ and the skewness matrix is generated considering the off-diagonal elements equal to zero, whereas the diagonal matrix is generated from $\operatorname{Uniform}(-0.5,0.5)$. We generate the $\operatorname{SUT}$ distribution with degrees of freedom coming from $\operatorname{Uniform}(2,10)$ [^2] The range is determined by estimating the degrees of freedom of the data on stock returns. . The estimators are then plugged into the ”Glasso” algorithm in (14) to obtain $\Omega$.
To evaluate the accuracy of the methods, we use the receiver operating curve ($\operatorname{ROC}$). We define the false positive rate and true positive rate for a fixed set of tuning parameters $\lambda_{G}$. Let $\hat{G}^{\lambda}_{G}=(V,\hat{E}^{\lambda_{G}})$ be an estimated graph using the proposed procedure and a given value of $\lambda_{G}$. We define the number of false positives for a given $\lambda_{G}$, $\operatorname{FP}(\lambda_{G})$, as the number of edges in $\hat{E}^{\lambda_{G}}$ but not in $E$ and the number of false negatives, $\operatorname{FN}(\lambda_{G})$, as the number of edges in $E$ but not in $\hat{E}^{\lambda_{G}}$. We consider the rate of the number of false negative and false positive:
@@
Graphically, we show averaged ROC curves for all the considered scenarios over 100 trials, plotted using the coordinates in (22). Fig. 1 shows numerical simulations of the average ROC curves of the elliptical skew-(S)KEPTIC estimators and the SKEPTIC estimators, which we employ to estimate $\Omega$ using the Glasso algorithm. The estimators exhibit similar performance, consequently yielding an optimal graph recovery in the simulation results, even when accounting for skewness and kurtosis components in the data, thus giving reliability to the elliptical skew-(S)KEPTIC estimators.
Table 2 presents the quantitative comparisons, showing that regardless of the data generation process or the estimator used, the structure of $\Omega$ can be effectively recovered. Incorporating a skewness component and accounting for heavier tails than those in the NPN distribution do not hinder the recovery of $\Omega$’s structure.
Table 2: Quantitative comparisons of the 4 estimators using different distributions to generate data, and compute the $\Omega$ matrix. The graphs are estimated using the Glasso algorithm. The values are in percentage.
SKEPTIC ($\hat{\rho}$) SKEPTIC ($\hat{\tau}$) SKEW-SKEPTIC ($\hat{\rho}$) SKEW-SKEPTIC ($\hat{\tau}$) $n$ $p$ FPR FNR FPR FNR FPR FNR FPR FNR N DATA 200 100 11.11 74.3 11.2 74.3 11.8 74.7 11.4 74.7 s.e (1.3) (3.5) (1.3) (3.5) (1.4) (3.5) (1.3) (3.5) 100 100 15.5 66.9 15.4 67.0 16.1 67.1 15.7 67.2 s.e (1.6) (2.4) (1.0) (2.4) (1.1) (2.4) (1.1) (2.4) 100 200 13.8 64.0 14.0 64.2 14.9 64.0 14.5 64.1 s.e (1.0) (2.4) (1.0) (2.4) (1.1) (2.4) (1.1) (2.4) CSN DATA 200 100 11.5 73.9 11.5 74.0 12.1 74.1 11.6 73.7 s.e (1.4) (3.5) (1.4) (3.5) (3.5) (3.5) (1.4) (3.5) 100 100 15.6 67.2 15.8 67.3 16.0 65.8 15.3 65.1 s.e (1.6) (3.4) (1.6) (3.4) (1.7) (3.5) (1.6) (3.5) 100 200 13.9 64.0 14.1 64.1 14.3 62.1 13.6 61.6 s.e (1.0) (2.4) (1.0) (2.4) (1.1) (2.5) (1.0) (2.5) SUT DATA 200 100 12.6 72.6 12.8 72.8 - - 14.9 61.0 s.e (1.5) (3.5) (1.5) (3.5) (-) (-) (1.7) (3.3) 100 100 16.8 64.4 17.0 64.5 - - 17.1 64.5 s.e (1.7) (3.3) (1.7) (3.3) (-) (-) (1.7) (3.3) 100 200 14.5 60.7 14.5 61.0 - - 14.9 61.0 s.e (1.0) (2.4) (1.0) (2.4) (-) (-) (1.1) (2.4)

### v2

@@ 4.2 The Glasso Algorithm
The penalty term, $\lambda_{G}$, is the parameter that controls the sparsity of the estimated matrix, i.e., the number of zeros in the precision matrix. Banerjee et al., (2008) shows that the shrinking procedure based on the tuning parameter $\lambda_{G}$ is directly proportional to the input covariance matrix. This means that greater values of $\lambda_{G}$ imply a greater presence of zeros in the estimated $\Omega$, therefore, a minor number of interconnections are unveiled.
4.3 The CLIME Algorithm
CLIME stands for constrained $\ell_{1}$-minimization for inverse matrix estimation. This algorithm, proposed by Cai et al., (2011), is very attractive for high-dimensional data since the estimator is obtained by combining vector solutions by solving a linear program. Let $\Omega$ denote a $p\times p$ precision matrix and $\{\hat{\Omega}_{1}\}$ be the solution set of the optimization problem:
$\text{min}\|\Omega\|_{1}\quad\text{subject to:}$ (12)
$\|\Sigma\Omega-I\|_{\infty}\leq\lambda_{C},$
where $\lambda_{C}$ is the tuning parameter. The final CLIME estimator of $\Omega$ is defined by the symmetrization of $\hat{\Omega}_{1}$:
$\hat{\Omega}=(\hat{\omega}_{ij}),\quad\text{where:}$ (13)
@@ 5 Theoretical properties
We investigate the theoretical properties of the elliptical skew-(S)KEPTIC estimators, establishing concentration inequalities in $\|\cdot\|_{\max}$ for our correlation matrix estimators, and the sparsistency of $\hat{\Omega}$ which describes the characteristic where any parameter that is originally zero is correctly estimated as zero, with a probability approaching one. For the rate of convergence of the correlation matrix estimator, we revisit the results of Liu et al., 2012a . For the corresponding precision matrix, we follow the framework introduced in Lam and Fan, (2009) for demonstrating the rate of convergence and the sparsistency of $\hat{\Omega}$. Our findings are based on the application of shrinkage procedures incorporating an $L_{1}$-penalty, with the three methods outlined in Section 4 adhering to this approach.
@@
Let $x_{1},...,x_{n}$ be $n$ observations from $\boldsymbol{X}\sim MS\mathcal{E}_{p,p}(\Omega^{-1},D_{\alpha},h^{(2p)},0,D_{\kappa};f)$, restricted to the case $h^{(p)}=\phi^{(p)}(u)$, and let $\hat{S}^{\rho,\alpha}_{ij}=4\sin(\frac{\pi}{6}A_{\hat{\rho,}_{ij}})\operatorname{I}(i\neq j)+\operatorname{I}(i=j)$ denote the elliptical skew-SKEPTIC, then with probability at least $1-\frac{1}{p}$ we have:
$\|\hat{S}^{\rho,\alpha}_{ij}-\Sigma^{0}_{ij}\|_{\max}\leq\sqrt{\frac{3(2\pi-3)^{2}}{(9-2\pi)^{2}}\frac{\log\left({p}\right)}{n}}.$ (15)
@@
Let $x_{1},...,x_{n}$ be $n$ observations from $\boldsymbol{X}\sim MS\mathcal{E}_{p,p}(\Omega^{-1},D_{\alpha},h^{(2p)},0,D_{\kappa};f)$ and let $\hat{S}^{\tau,\alpha}_{ij}=2\sin(\frac{\pi}{2}A_{\hat{\tau}_{ij}})\operatorname{I}(i\neq j)+\operatorname{I}(i=j)$ denote the elliptical skew-(S)KEPTIC estimators, then with probability at least $1-\frac{1}{p}$ we have:
$\|\hat{S}^{\tau,\alpha}_{ij}-\Sigma^{0}_{ij}\|_{\max}\leq\sqrt{{6\pi^{2}}\frac{\log\left({p}\right)}{n}}.$ (16)
The proofs of Theorems 5.1 and 5.2 are in Appendix A. The elliptical skew-(S)KEPTIC estimators are consistent, so when the sample size $n$ increases, the estimator converges to the true correlation matrix.
Let $\mathcal{S}=\{(i,j):\omega^{0}_{ij}\neq 0\}$ be the support set of off-diagonal nonzero entries, with cardinality $|\mathcal{S}|=s+p$, let $p_{\lambda}(\cdot)$ be a penalty function, and define
$a=\max_{(i,j)\in\mathcal{S}}p^{\prime}_{\lambda}(|\ell_{ij}|),\quad b=\max_{(i,j)\in\mathcal{S}}p^{\prime\prime}_{\lambda}(|\ell_{ij}|).$
where $\ell_{ij}$ is the $(i,j)$ element of the $L$ matrix that stems from $\Omega=L^{\top}ML$. In this work, we consider shrinkage procedures based on $L_{1}$-penalty, which means $a=\lambda$ and $b=0$. We impose the following regularity conditions:
  • (a)
    The eigenvalues of $\Sigma_{0}$ are bounded: $0<\tau_{1}<\lambda_{\min}(\Sigma_{0})\leq\lambda_{\max}(\Sigma_{0})<\tau_{2}<\infty$.
  • (b)
    Assume $a=O\left(\sqrt{\frac{(1+p/(s+1))\log p}{n}}\right)$, $b=o(1)$, and $\min_{(i,j)\in\mathcal{S}}|\omega^{0}_{ij}|/\lambda\to\infty$.
  • (c)
    The penalty is singular at the origin: $\lim_{t\downarrow 0}p_{\lambda}(t)/(\lambda t)=k>0$.
  • (d)
    For some constants $C,D>0$, when $\theta_{1},\theta_{2}<C\lambda$,
    $|p^{\prime\prime}_{\lambda}(\theta_{1})-p^{\prime\prime}_{\lambda}(\theta_{2})|\leq D|\theta_{1}-\theta_{2}|.$
@@
Suppose that conditions (a), (b), (c), (d) hold. If $\frac{\log p}{n}=O(\lambda^{2})$, then there exists a local minimizer for the maximum likelihood estimator for $\widehat{{M}}$ and an estimate of $\widehat{{L}}$ that satisfy:
$\|\widehat{{M}}-{M}\|_{\max}=O_{P}\left(\sqrt{\frac{\log p}{n}}\right),\quad\|\widehat{{L}}-{L}\|_{\max}=O_{P}\left(\sqrt{\frac{\log p}{n}}\right),$
which implies:
$\|{\Omega}-\Omega\|_{\max}=O_{P}\left(s\sqrt{\frac{\log p}{n}}\right).$
The proof of Theorem 5.3 is in Appendix A. We know from Lemma 1 in Zareifard et al., (2016) that $L_{ij}=0\Leftrightarrow\Omega_{ij}=0$, so the matrices $L$ and $\Omega$ share the same sparsity patter. For this reason, we provide sparsistency directly on $\Omega$.
@@
(Sparsistency) Under the conditions (a), (b), (c), (d) for a local minimizer based on $L_{1}$-penalty that satisfies $\|\widehat{\Omega}-\Omega\|_{\max}^{2}=O_{P}\left\{s\log p/n\right\}$ and $\|\widehat{\Omega}-\Omega\|^{2}=O_{P}\left(\eta_{n}\right)$ for a sequence of $\eta_{n}\rightarrow 0$, where $\eta_{n}\equiv s\sqrt{\frac{\log p}{n}}$. If $\log p/n+\eta_{n}=O\left(\lambda_{n1}^{2}\right)$, then with probability tending to $1,\widehat{\omega}_{ij}=0$ for all $(i,j)\in S^{c}$.
The proof of Theorem 5.4 is in Appendix A.
@@ 6.1 Numerical Simulations
We perform numerical simulations by generating synthetic data from the normal distribution, the closed skew-normal distribution, and the unified skew-$t$ distribution. For each dataset, we compute both the SKEPTIC and the skew-(S)KEPTIC estimators, and we evaluate the graph recovery performance of the proposed methods.
The implementation follows these steps. Using the R package huge (Jiang et al., (2021)), we generate the precision matrix $\Omega=(\Sigma^{0})^{-1}$ using the ”random” structure, in which we randomly generate pairs of conditional dependencies in the off-diagonal elements. We sample $n$ observations from $p$-variate normal, CSN and SUT distribution. We set the location parameter equal to zero, and $\Sigma^{0}=\Omega^{-1}$.
We consider the following cases: $p\in\{100,250,300\}$, and $n\in\{200,1000,500\}$. We generate synthetic data assuming $f(\boldsymbol{X})\sim N(0,\Sigma^{0})$, $f(\boldsymbol{X})\sim\operatorname{CSN}(0,\Sigma^{0},D_{\alpha}L,0,D_{\kappa})$, and $f(\boldsymbol{X})\sim\operatorname{SUT}(0,\Sigma^{0},D_{\alpha}L,0,D_{\kappa})\equiv\frac{W}{\sqrt{V_{0}}}$, where $W$ is distributed according to a closed skew-normal distribution, and $V_{0}\sim Gamma(\nu/2,\nu/2)$ as in Proposition 2.5. We set $\nu=3$ to emphasize the heavy tails in the distribution. To generate the data from the closed skew-normal distribution, we rely on the stochastic representation in (7). Moreover, we generate the skewness parameters from $\alpha=0.4\cdot\operatorname{sign}(u-0.5)\cdot\sqrt{|u-1|},\quad\text{where }u\sim\mathcal{U}(0,1)$. These parameters are encoded in the diagonal of the matrix $D_{\alpha}$.
After the data generation, we compute the skew-(S)KEPTIC and the SKEPTIC estimators, which are then plugged into the glasso algorithm in (11) to obtain $\Omega$. The glasso algorithm depends on a tuning parameter $\lambda_{G}$ whose the choice is performed applying the StARS method in Liu et al., (2010), and when the value is found for each dataset we consider a set of 30 tuning parameters in the range $(\lambda_{G}-0.1,\lambda_{G}+0.1)$.
To evaluate the accuracy of the methods, we use the receiver operating curve ($\operatorname{ROC}$). We define the false positive rate and true positive rate for a fixed set of tuning parameters $\lambda_{G}$. When the tuning parameter is determined we estimate the graph: $\hat{G}^{\lambda_{G}}=(V,\hat{E}^{\lambda_{G}})$, to evaluate it, we define the number of false positives for a given $\lambda_{G}$, $\operatorname{FP}(\lambda_{G})$, as the number of edges in $\hat{E}^{\lambda_{G}}$ but not in $E$ and the number of false negatives, $\operatorname{FN}(\lambda_{G})$, as the number of edges in $E$ but not in $\hat{E}^{\lambda_{G}}$. We consider the rate of the number of false negatives and false positives:
@@
Graphically, we show averaged ROC curves for all the considered scenarios over 100 trials, plotted using the coordinates in (18). Fig. 1 shows that the skew-(S)KEPTIC estimators generally outperform the Pearson correlation approach in non-Gaussian and improve upon the standard SKEPTIC estimator. Adjusting for skewness, also for excess kurtosis in the case of the SUT distribution, enhances the estimation of $\Omega$, especially in settings where the distribution presents heavy tails. In some cases, this leads to better recovery compared to the original SKEPTIC estimator, which has been shown to be optimal in Liu et al., 2012a . Table 2 provides a comprehensive overview of the simulation results, presenting quantitative comparisons of various estimators. The average AUC values of the skew-(S)KEPTIC estimators highlight strong performance in graph recovery under non-Gaussian conditions while still maintaining reliability when applied to normally distributed data. These findings suggest that considering estimators that include the skewness and accommodate for heavier tails, extending the nonparanormal distribution, do not hinder the recovery of the $\Omega$ structure. On the contrary, they may offer a practical advantage due to the robustness across different data distributions.